OpenStudy (texmechanica):

Complete the factorization. a2 + 3a – 28 = (a – 4)(a +

OpenStudy (ilikeramen):

Hi Fweind

OpenStudy (ilikeramen):

Whut u think it is :D

OpenStudy (texmechanica):

The 2 after the A is an exponent by the way.

OpenStudy (ilikeramen):

a+6

OpenStudy (texmechanica):

Is that the answer?

OpenStudy (texmechanica):

Is the 6 after the a, or is a+6 after the a at the end?

OpenStudy (anonymous):

What two numbers can you multiply together to make 28 and add up to 4?

OpenStudy (texmechanica):

Can you just show me how to factor?

OpenStudy (anonymous):

Break 28 up into its prime factors and you should see it. I won't solve it for you because you need to know how to do it yourself. I'll help you.

OpenStudy (anonymous):

28 factors into 2 * 14. What are the factors of 14?

OpenStudy (anonymous):

What two numbers can you multiply together to make 14 besides 1 and 14?

OpenStudy (texmechanica):

2 and 7.

OpenStudy (anonymous):

if u wanna factor the only thing u should do is to say: ____ + ____ = 3 ____ * _____ = -28 now u can say -4 and 7 lets check it : 7 - 4 = 3 7 * -4 = -28 :))) i hope this is helpful have a nice day

OpenStudy (anonymous):

If you break the last term up into its prime factors you should be able to find a combination of those factors that when multiplied together equal the last term and when added together equal the middle term. 2 * 2 * 7 = 28 4 * 7 = 28 7-4=3

OpenStudy (texmechanica):

Is 3 the answer?

OpenStudy (anonymous):

no the answer is 7